What will be the temperature change of 0.5 kg of water(c= 4200J/kg K) that loses 21000J.


1
Expert's answer
2021-06-09T16:14:31-0400
"Q=mc\Delta T,""\Delta T=\dfrac{Q}{mc},""\Delta T=\dfrac{21000\ J}{0.5\ kg\cdot4200\ \dfrac{J}{kg\cdot \!^{\circ}C}}=10^{\circ}C."